Functions f , g : R → R are defined, respectively, by f (x) = x2 + 3x + 1, g (x) = 2x – 3,
find   g o g 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
f(x)=x^2+3x+1
g(x)=2x–3

gog = g(g(x)) = g(2x - 3)

= 2(2x - 3)–3
= 4x–6–3
= 4x - 9


Thanks